Shift report 20000806 08:00-16:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Sun Aug 06 2000 - 18:14:13 EDT

  • Next message: Konstantin Olchanski: "New High Voltage monitoring web page is now available"

    ------------------------------------------------------------
    Date: Sun Aug  6 18:14:12 2000, Shift: 08:00-16:00
    Supervisor: R. D., with: T. K.
    Log book#4 pages 11 - 14
    
    Shift Summary:
    ------------------------------------------------------------
    At 14:20 we finally got a good store at gamma 70. It is still
    going.
    
    We took runs 2072-2074 all with D5 on at 1/4 of full field.
    
    
    Once again the vertex appears to be shifted, we informed MCR but were told that for this store they would not change the cogging
    We also reported that the rate of collisions was 1/2 per second.
    They also said that they will steer beam at Phenix during next shift, and then
    they may look at our rate.
    
    H1 and C1 remain "incomunicados" because of HV mainframe problem
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Sun Aug 06 2000 - 18:17:23 EDT